# On the analytic continuation of Dirichlet series with missing digits

### v1

@@ 1. Introduction
Consider the Cantor set of the numbers in $[0,1]$ whose (infinite) $b$-radix representation (or one of them, if there are two) uses only admissible digits (it $N=1$ it is reduced to a singleton and if $N=b$ it is the un-modified interval itself). It is well-known that there is a canonical probability measure $\mu$ on this set, singular to Lebesgue measure if $N<b$, whose moment generating function $\int_{[0,1]}\mathsf{e}^{tx}\mathrm{d}\mskip-1.0mu\mu(x)$ is the infinite product (which is an entire function):
@@ 2. Representation by geometrically convergent series
A general basic result (Jensen lemma, [17, Chap. 4]), underlying the notion of abscissa of convergence, says that as a consequence, the Dirichlet series (1) defining $K(s)$ converges if and only if $\Re(s)>s_{0}$ ([15], see also [16] for more generality, and [4] for a still wider perspective).

### v2

@@ 1. Introduction
Consider the Cantor set of the numbers in $[0,1]$ whose (infinite) $b$-radix representation (or one of them, if there are two) uses only admissible digits (if $N=1$ it is reduced to a singleton and if $N=b$ it is the un-modified interval itself). It is well-known that there is a canonical probability measure $\mu$ on this set, singular to Lebesgue measure if $N<b$, whose moment generating function $\int_{[0,1]}\mathsf{e}^{tx}\mathrm{d}\mskip-1.0mu\mu(x)$ is the infinite product (which is an entire function):
@@ 2. Representation by geometrically convergent series
A general basic result (Jensen lemma, based upon a partial summation, see [17, Chap. 4]), which underlies the notion of abscissa of convergence, says that as a consequence, the Dirichlet series (1) defining $K(s)$ diverges whenever $\Re(s)<s_{0}$. The abscissa of convergence is thus $s_{0}$ ([15]; see also [16] for more generality, and [4] for a still wider perspective). Let us mention this complement:
Proposition 2.1.
For any $s=s_{0}+it$, $t\in\mathbb{R}$, the Dirichlet series $K(s)$ diverges.
Proof.
Suppose to the contrary that it converges for some $s=s_{0}+it$. Let $n_{1}<n_{2}<\dots<n_{q}<\dots$ be the positive admissible integers. Let $r_{p}=\sum_{q=p}^{\infty}n_{q}^{-s}=o(1)$. We do a summation by parts:
$S_{p}\coloneq\sum_{q=1}^{p}n_{q}^{-s_{0}}=\sum_{q=1}^{p}(r_{q}-r_{q+1})n_{q}^{it}=r_{1}n_{1}^{it}+\sum_{q=2}^{p}r_{q}\bigl(n_{q}^{it}-n_{q-1}^{it}\bigr)-r_{p+1}n_{p}^{it}.$
From $n_{q}^{it}-n_{q-1}^{it}=\int_{n_{q-1}}^{n_{q}}itx^{it-1}\mathrm{d}x$, or from $|\mathsf{e}^{i\theta}-1|\leq|\theta|$, we get $\big|n_{q}^{it}-n_{q-1}^{it}\bigr|\leq|t|\log\frac{n_{q}}{n_{q-1}}$. The summation on the right-hand side, restricted to the indices $q$ such that $p_{0}<q\leq p$ with some fixed $p_{0}$ is thus bounded by $|t|(\sup_{q>p_{0}}|r_{q}|)\log(n_{p}n_{p_{0}}^{-1})$. Bounding trivially the finitely many other terms, it follows:
$\lim_{p\to\infty}\frac{S_{p}}{\log n_{p}}=0.$
But this is wrong. To see it, take $\ell\geq 1$ and $p_{\ell}=N_{1}+N_{1}N+\dots+N_{1}N^{\ell-1}$ so that $n_{p_{\ell}}$ is the largest admissible integer which is $b^{\ell}$. Thus $n_{p_{l}}<b^{\ell}$. And $S_{p_{\ell}}>N_{1}b^{-s_{0}}+N_{1}Nb^{-2s_{0}}+N_{1}N^{2}b^{-3s_{0}}+\dots+N_{1}N^{\ell-1}b^{-\ell s_{0}}=\ell N_{1}N^{-1}$. Hence, $S_{p_{\ell}}$ is not $o(\log n_{p_{\ell}})$. ∎